OpenStudy (anonymous):

How do you "Calculate the average speed of the ball on the following intervals" given the graph of the height of a ball thrown into the air

OpenStudy (anonymous):

\[s = \frac{ \Delta x }{ \Delta t }\]

OpenStudy (anonymous):

y=1/2gt^2+v0t + y0

OpenStudy (anonymous):

So calculate the speed of at each point of the interval, meaning plug in t. Then the rate of change in distance at the certain time intervals.

OpenStudy (anonymous):

Once you have change in distance and time you can calculate average speed.

OpenStudy (anonymous):

or if you your have not T as use this formula: v^2-v0^2=2ah

OpenStudy (anonymous):

this graph is given

OpenStudy (anonymous):

then it says for [0,.5] but 23/.5 isn't correct

OpenStudy (anonymous):

what's qu(5) ? and what kind of graph is this? position vs time? velocity vs time???

OpenStudy (anonymous):

The vertical units are meters and the horizontal units seconds.

OpenStudy (anonymous):

than how is 23/0.5 not correct? unless the height at t = 0.5s isn't 23m